# Real-time Scattering in $\mathbf{\phi^{4}}$ Theory using Matrix Product States

### v1

@@ III Results
Finite-entanglement scaling.— Using uMPS with bond dimensions $D=\{48,64,96\}$, we performed a coarse-grained scan of the mass-squared parameter $r\equiv\mu^{2}$ over a broad interval and identified an approximate critical region around $r\simeq-0.32$. We then carried out a refined scan in the window $r\in[-0.325,-0.310]$ using the same FES procedure. For each pair $(r,D)$ we extracted the correlation length $\xi_{D}(r)$ from the subleading eigenvalue of the uMPS transfer matrix and the bipartite entropy $S_{D}(r)$ from the Schmidt spectrum, and fitted $S_{D}(r)$ as a function of $\log\xi_{D}(r)$ to obtain the slope $a(r)$ and effective central charge $c_{\mathrm{eff}}(r)=6\,a(r)$, as described in the Methods.
(a) $c_{\text{eff}}$ vs $\mu^{2}$.
(b) $\langle\phi\rangle$ across the phase transition.
(c) $\mathcal{S}$ vs $\log\xi$.
(d) $\xi$ vs $D$
@@
Figure 1: Finite-entanglement scaling near the critical point of the $\phi^{4}$ theory at $\lambda=0.8$. Ground-state uMPS data with finite bond dimension $D$ are used to extract the effective central charge $c_{\mathrm{eff}}(r)$, a local order parameter, and the entropy–correlation-length relation near $r_{c}=\mu_{c}^{2}$. The peak in $c_{\mathrm{eff}}(r)$ and the linear dependence $\mathcal{S}\simeq\frac{c}{6}\log\xi_{D}+\mathrm{const.}$ close to $r_{c}$ yield the bracket $\mu_{c}^{2}\in[-0.3190,-0.3185]$ and confirm Ising universality.
@@
Across the sweep, $c_{\mathrm{eff}}(r)$ grows rapidly as $r$ approaches $r\simeq-0.319$ and exhibits a clear maximum within our grid:
$c_{\mathrm{eff}}(-0.3190)$ $=0.118\pm 0.022,$
$c_{\mathrm{eff}}(-0.3185)$ $=0.191\pm 0.102,$
$c_{\mathrm{eff}}(-0.3180)$ $=0.063\pm 0.015.$ (7)
@@
$r_{c}\in[-0.3190,-0.3185],$ (8)
with a central estimate near $r_{c}\approx-0.31875$. Given the modest entanglement range $D\in[48,96]$ and the relatively coarse sampling in $r$, the peak height underestimates the Ising value $c=1/2$, as expected from finite-$D$ truncation, but the trend of $c_{\mathrm{eff}}(r)$ is fully consistent with Ising universality.
For the subsequent scattering study we use this FES map to select representative couplings. In particular, we choose a symmetric-phase point $r=+0.2$, a near-critical point close to the bracketed $r_{c}$, and a weakly broken point r=0.2r=-0.2 within the spontaneously broken phase. These choices ensure that the scattering simulations sample distinct regions of the phase diagram while remaining quantitatively anchored to the FES estimate of $r_{c}$. A more deeply broken point at r=0.5r=-0.5 is also included to illustrate the strongly gapped regime.
Scattering in the $\phi^{4}$ theory.— Before studying real-time dynamics, we characterize the excitation spectrum through the single-particle and multi-particle bands shown in Fig. 2. In the symmetric phase at $\mu^{2}=+0.2$ [Fig. 2(a)], the system is gapped: the lowest excitation at $p=0$ has energy $dE\approx 1.2$, and there is a clear gap between the single-particle band and the multi-particle continuum, which onsets around $dE\approx 2.4$. This is characteristic of a stable symmetric phase, with a vacuum at $\phi=0$ and standard massive bosonic excitations.
@@
Very close to the critical point at $\mu^{2}_{c}\approx-0.31875$ [Fig. 2(b)] the spectrum becomes gapless. The mass of the fundamental particle vanishes ($m=0$), and the lowest band starts at $dE\approx 0$, while the multi-particle thresholds (at $2m$, $3m$, etc.) collapse onto one another. The spectrum is consistent with an underlying conformal field theory: the dispersion near $p=0$ is approximately linear, $E\propto|p|$, and the stacked structure of the low-lying bands suggests a tower of stable or long-lived bound states at this coupling. In the spontaneously broken phase [Figs. 2(c), 2(d)], the system is gapped again and a single-particle band re-emerges with a nonzero mass $m_{s}$. As the system moves away from criticality into the deeply broken regime (e.g., $\mu^{2}=-0.5$), this mass increases, providing a spectral signature of mass generation across the phase diagram.
To probe the nonperturbative dynamics we simulate real-time two-particle collisions across these regimes, as shown in Fig. 3. In the gapped phases, panels [3(a)], [3(c)], and [3(d)] display the characteristic diagonal “X" pattern in $\langle\phi_{n}(t)\rangle$, indicative of two localized wave packets approaching, interacting, and separating again as outgoing quasiparticles. The overall structure confirms the stable propagation of massive excitations in both the symmetric and spontaneously broken phases.
@@
$\Delta t_{\mathrm{sym}}(E_{0})$ $\simeq-180.718.$ (10)
signaling a substantial cross section for particle production and other inelastic channels, as expected for a nonintegrable theory with a fundamental boson. By contrast, in the weakly broken phase [Fig. 3(c), $\mu^{2}=-0.2$] the scattering is almost perfectly elastic,
$P_{11\rightarrow 11}(E_{0})$ $\simeq 0.998,$ (11)
$\Delta t_{\mathrm{br}}(E_{0})$ $\simeq-269.720,$ (12)
with the incoming wave packets essentially re-emerging as outgoing packets with the same particle content and minimal radiation. The deeply broken point [Fig. 3(d), $\mu^{2}=-0.5$] exhibits similar qualitatively elastic behavior, $P_{11\rightarrow 11}\simeq 1$ and $\tau\simeq-177.781$, with even more clearly separated outgoing wave fronts due to the larger gap.
At the critical coupling $\mu_{c}^{2}=-0.31875$ [Fig. 3(b)], the sandwich evolution ceases to produce a well-defined scattering event. Instead of the characteristic “X”-shaped collision pattern, $\langle\phi_{n}(t)\rangle$ develops a long-wavelength drift across the entire window, and no clean separation into incoming and outgoing wave packets is observed. This is not a numerical artifact or an instability of TDVP/uMPS, but a direct consequence of criticality: the sandwich construction assumes a finite correlation length $\ell$ that separates a central scattering region from asymptotic bulk domains. At the critical point the exact ground state is gapless ($m=0$) and scale invariant, with a diverging correlation length ($\ell\to\infty$), so any finite-window implementation of this protocol inevitably violates that assumption and fails to reach asymptotic scattering states. In this sense, the breakdown of the sandwich geometry at $\mu^{2}=\mu_{c}^{2}$ reflects the underlying physics of an infinitely correlated critical state rather than a flaw of the method, and it provides a dynamical signature of the quantum critical point that is complementary to the static FES diagnostics discussed above. Closer and closer to $\mu_{c}^{2}$ the required window sizes and evolution times grow with $\ell$, so in practice the protocol becomes rapidly more expensive and ultimately impractical as the critical point is approached.
@@ IV Summary and Discussion
We have extended TDVP-based uniform matrix product state techniques, previously applied to Ising field theory [1], to the interacting $\phi^{4}$ quantum field theory in $(1+1)$ dimensions. A finite-entanglement scaling analysis at $\lambda=0.8$ yields a controlled estimate of the critical mass-squared, $\mu_{c}^{2}\in[-0.3190,-0.3185]$, consistent with Ising universality and with the expected emergence of a gapless, conformal regime. This provides a quantitative map of the symmetric, near-critical, weakly broken, and deeply broken phases, anchoring the subsequent real-time simulations to a data-driven bracket of the critical point rather than to ad hoc parameter choices.
Using these uMPS ground states as asymptotic vacua, we simulated two-particle collisions in a sandwich geometry across the phase diagram and extracted both the elastic scattering probability $P_{11\to 11}(E)$ and the Wigner time delay $\Delta t(E)$. In the symmetric phase the collisions are strongly inelastic, with $P_{11\to 11}\approx 0.63$, reflecting the sizable phase space for particle production in a nonintegrable theory with a fundamental boson. In the spontaneously broken phase the same protocol yields almost perfectly elastic collisions, $P_{11\to 11}\approx 0.998$, indicating that the excitations behave as stable, weakly radiating quasiparticles. As the system is tuned deeper into the broken regime, the increasing gap sharpens this behavior and further suppresses inelastic channels.
At the critical coupling $\mu^{2}=\mu_{c}^{2}$ the sandwich evolution fails to produce a clean scattering event: instead of a well-defined “X” pattern in $\langle\phi_{n}(t)\rangle$, we observe a long-wavelength drift across the entire window and no clear separation into incoming and outgoing packets. This breakdown is a direct manifestation of criticality: the scattering construction relies on a finite correlation length to isolate a central interaction region from asymptotic bulk domains, an assumption that is invalidated when the mass gap closes and $\xi\to\infty$. The failure of the sandwich protocol at $\mu^{2}_{c}$ thus provides a dynamical signature of the quantum critical point, complementary to the static FES diagnostics.
@@ Acknowledgments
W.C. gratefully acknowledges the support provided by Jens Eisert, including financial assistance from Freie Universität Berlin through a research fellowship funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the Research Unit FOR 2724 “Thermische Maschinen in der Quantenwelt" (Fonds 0420611402).

### v2

@@ III Results
III.1 Validation against Known Limits
Before investigating finite entanglement scaling or real-time scattering, we validate our uMPS framework and Hamiltonian discretization by reproducing known regimes of the (1+1)(1+1)D $\phi^{4}$ theory: the free-particle dispersion, the location of the critical point, the verification of the $\beta$-parameter, and the characterization of the broken phase.
III.1.1 Single-Particle Dispersion
To verify that our ansatz correctly captures the kinetic terms and momentum eigenstates, we computed the excitation spectrum in the symmetric phase. Fig. 3 compares the MPS-derived excitation energies $E(p)$ against the analytic lattice dispersion relation for a massive boson,
$E^{2}(p)=\mu^{2}+4\sin^{2}(p/2).$ (19)
The simulation data shows a good agreement with the analytic curve (dashed black line) across the Brillouin zone. This confirms that our bond dimension $D=32$ and our Hilbert space truncation $d=8$ are sufficient to resolve the single-particle band without introducing significant lattice artifacts or finite-entanglement distortions.
We have taken here $\mu^{2}=0.2$.
Figure 3: Verification of the single-particle dispersion relation in the free theory ($\lambda=0,\;\mu^{2}=0.2$). The MPS excitation spectra (for different bond dimensions and Hilbert-space truncations) match closely the analytic lattice dispersion (black dashed line).
III.1.2 Critical Point Location
We locate the quantum critical point (QCP) by monitoring the mass gap $\Delta$ as a function of the quartic coupling $\lambda$ at fixed $\mu^{2}=-1.0$. We have taken $D=32$ and $d=12$.
The $\phi^{4}$ transition in (1+1) dimensions belongs to the 2D Ising universality class. In this class, the correlation length $\xi$ diverges at the critical point as
$\xi\propto\frac{1}{|\lambda-\lambda_{c}|^{\nu}},$ (20)
where $\nu=1$ for this class. The mass (energy) gap $\Delta$ is inversely proportional to the correlation length and so
$\Delta\sim|\lambda-\lambda_{c}|.$ (21)
As shown in Fig. 4, the system exhibits a clear “V-shape” gap closing, minimizing at $\lambda_{c}\approx 3.79$. This closing separates the symmetric phase (single-well effective potential) from the broken phase (double-well potential), demonstrating the method’s ability to capture the long-range correlations associated with the phase transition.
Figure 4: Detection of the phase transition via gap closing. The energy gap (blue line) vanishes at a critical coupling $\lambda_{c}\approx 3.79$ (for fixed $\mu^{2}=-1.0$), marking the transition from the symmetry-broken phase to the symmetric phase.
III.1.3 The ‘Sugihara Ratio’ Test
As a further check for the accuracy of our uMPS simulations, we look at the work of T. Sugihara [37].
Figure 5: The divergent diagram in (1+1)D $\phi^{4}$ model.
In the $\phi^{4}$ model, there is only one diagram that diverges in the continuum limit $a\rightarrow 0$ (see Fig. 5). We can renormalize it by redefining the mass parameter as follows:
$\mu^{2}=\mu^{2}_{\mathrm{ren}}-\delta\mu^{2},$ (22)
where $\delta\mu^{2}$ is the tadpole counter-term that cancels the divergence and is defined as:
$\delta\mu^{2}=\frac{\lambda}{2\pi\sqrt{\mu_{\mathrm{ren}}^{2}+4}}K\left(\frac{4}{\mu^{2}_{\mathrm{ren}}+4}\right),$ (23)
with $K(x)$ is the complete elliptic integral of the first kind.
Sugihara confirms, using density matrix renormalization group (DMRG) calculations, that the ratio
$R_{c}=\frac{\lambda}{\mu_{\mathrm{ren}}^{2}}\bigg|_{\mathrm{critical~point}}=59.89\pm 0.01.$ (24)
Table 1 presents a convergence study comparing our simulated results, $R_{uMPS}$, against this theoretical target across various parameters. Because any numerical simulation using Matrix Product States is constrained by bond dimension $D$ and the local Hilbert space truncation $d$ we expect a measurable deviation from $R_{c}$.
For example, at a fixed coupling of $\lambda=3.0$, increasing the bond dimension from $D=32$ to $D=64$ reduces the deviation from $4.43\%$ to $3.32\%$, demonstrating that our results systematically approach the continuum benchmark as computational power increases. The larger $11.30\%$ deviation observed at $\lambda=3.79$ highlights the increased numerical challenge at higher interaction strength. Ultimately, these results confirm that our uMPS framework remains well within an acceptable range of theoretical expectations, providing a stable foundation for our subsequent real-time scattering analysis.
$D$ $d$ $\lambda$ $\mu^{2}$ $\mu^{2}_{\mathrm{ren}}$ $R_{\mathrm{uMPS}}$ Deviation from $R_{c}$
$32$ $18$ $3.79$ 1.0-1.0 $0.05686$ $66.66$ $11.30\,\%$
$64$ $12$ $3.0$ 0.792-0.792 $0.05181$ $57.9$ $3.32\,\%$
$32$ $12$ $3.0$ 0.809-0.809 $0.04801$ $62.48$ $4.32\,\%$
$32$ 88 $0.8$ 0.25945-0.25945 $0.01243$ $64.36$ $7.46\,\%$
Table 1: Comparison between the ratio $R_{\mathrm{uMPS}}$ calculated from uMPS simulations and $R_{c}$ from Ref. [37]. We notice some deviation that is dependent on the bond dimension and the Hilbert space truncation, however, it lies within an acceptable range.
III.1.4 Symmetry Breaking and Quantum Fluctuations
In the spontaneously broken phase, quantum fluctuations are expected to suppress the order parameter $\langle\phi\rangle$ relative to the classical mean-field prediction. We tested this by simulating the system with a small pinning field ($h=0.1$) to select a vacuum. Here we have taken $\mu^{2}=-4.0$, $D=32$, and $d=24$.
Fig. 6 (left panel) shows a convergence to the classical value of $\sqrt{2|\mu^{2}|}$ at high lambdas although it drifts away at low lambdas. This is because at low lambda the true vacuum moves very far out requiring a bigger Hilbert space size $d$ to capture the precise mass gap. The right panel exhibits similar behavior, it approaches the classical limit, $1/\lambda$, for higher values of lambda.
Figure 6: Characterization of the broken phase with a pinning field $h=0.1$. (Left) The measured mass gap (blue) compared to the classical limit. (Right) The vacuum expectation value $\langle\phi\rangle$ (green) is suppressed relative to the classical prediction (red dashed), highlighting the method’s ability to capture quantum fluctuations.
III.2 Finite-entanglement scaling
@@
Using uMPS with $\lambda=0.8$ and bond dimensions $D=\{36,48,64\}$, we performed a coarse-grained scan of the mass-squared parameter $r\equiv\mu^{2}$ over a broad interval and identified an approximate critical region around $r\simeq-0.259$. We then carried out a refined scan in the window $r\in[-0.261,-0.258]$ using the same FES procedure. For each pair $(r,D)$ we extracted the correlation length $\xi_{D}(r)$ from the subleading eigenvalue of the uMPS transfer matrix and the bipartite entropy $S_{D}(r)$ from the Schmidt spectrum, and fitted $S_{D}(r)$ as a function of $\log\xi_{D}(r)$ to obtain the slope $a(r)$ and effective central charge $c_{\mathrm{eff}}(r)=6\,a(r)$, as described in the Methods.
(a) $c_{\mathrm{eff}}$ vs $\mu^{2}$.
(b) $\mathcal{S}$ vs $\log\xi$.
@@
Figure 7: Finite-entanglement scaling near the critical point of the $\phi^{4}$ theory at $\lambda=0.8$. Ground-state uMPS data with finite bond dimension $D$ are used to extract the effective central charge $c_{\mathrm{eff}}(r)$, and the entropy—correlation-length relation near $r_{c}=\mu_{c}^{2}$. The intersection with $c_{\mathrm{eff}}=0.5$ and the linear dependence $\mathcal{S}\simeq\frac{c}{6}\log\xi_{D}+\mathrm{const.}$ close to $r_{c}$ yield the bracket $\mu_{c}^{2}\in]-0.2595,-0.2594[$ and verify Ising universality.
@@
Across the sweep, $c_{\mathrm{eff}}(r)$ approaches $0.5$ as $r\rightarrow-0.2596$ and exhibits the following behavior in our grid [Fig. 7(b)]:
$c_{\mathrm{eff}}(-0.2596)$ $=0.303,$
$c_{\mathrm{eff}}(-0.2595)$ $=0.438,$
$c_{\mathrm{eff}}(-0.2594)$ $=0.542,$
$c_{\mathrm{eff}}(-0.2593)$ $=0.690,$
@@
$r_{c}\in]-0.2595,-0.2594[.$ (25)
We will take $r_{c}\approx-0.25945$.
For the subsequent scattering study we use this FES map to select representative couplings. In particular, we choose a symmetric-phase point $r=+0.2$, a near-critical point close to the bracketed $r_{c}$, and a weakly broken point r=0.1r=-0.1 within the spontaneously broken phase. These choices ensure that the scattering simulations sample distinct regions of the phase diagram while remaining quantitatively anchored to the FES estimate of $r_{c}$. A more deeply broken point at r=0.5r=-0.5 is also included to illustrate the strongly gapped regime.
III.3 Scattering in the $\mathbf{\phi^{4}}$ theory
Before studying real-time dynamics, we characterize the excitation spectrum through the single-particle and multi-particle bands shown in Fig. 8. In the symmetric phase at $\mu^{2}=+0.2$ [Fig. 8(a)], the system is gapped: the lowest excitation at $p=0$ has energy $dE\approx 0.6$, and there is a clear gap between the single-particle band and the multi-particle continuum, which onsets around $dE\approx 1.2$. This is characteristic of a stable symmetric phase, with a vacuum at $\phi=0$ and standard massive bosonic excitations.
@@ III.3 Scattering in the $\mathbf{\phi^{4}}$ theory
Very close to the critical point at $\mu^{2}_{c}\approx-0.2594$ [Fig. 8(b)] the spectrum becomes gapless. The mass of the fundamental particle vanishes ($m=0$), and the lowest band starts at $dE\approx 0$, while the multi-particle thresholds (at $2m$, $3m$, etc.) collapse onto one another. The spectrum is consistent with an underlying conformal field theory: the dispersion near $p=0$ is approximately linear, $E\propto|p|$, and the stacked structure of the low-lying bands suggests a tower of stable or long-lived bound states at this coupling.
In the spontaneously broken phase [Figs. 8(c), 8(d)], the system is gapped again and a single-particle band re-emerges with a nonzero mass $m_{s}$. As the system moves away from criticality into the deeply broken regime (e.g., $\mu^{2}=-0.5$), this mass increases, providing a spectral signature of mass generation across the phase diagram.
To probe the nonperturbative dynamics we simulate real-time two-particle collisions across these regimes, as shown in Fig. 9. In the gapped phases, panels [9(a)], [9(c)], and [9(d)] display the characteristic diagonal “X” pattern in $\langle\phi_{n}(t)\rangle$, indicative of two localized wave packets approaching, interacting, and separating again as outgoing quasiparticles. The overall structure confirms the stable propagation of massive excitations in both the symmetric and spontaneously broken phases.
@@
$\Delta t_{\mathrm{sym}}(E_{0})$ $\simeq-158.$ (27)
signaling a substantial cross section for particle production and other inelastic channels, as expected for a nonintegrable theory with a fundamental boson. By contrast, in the weakly broken phase [Fig. 9(c), $\mu^{2}=-0.1$] the scattering is almost perfectly elastic,
$P_{11\rightarrow 11}(E_{0})$ $\simeq 1,$ (28)
$\Delta t_{\mathrm{br1}}(E_{0})$ $\simeq-108.2158,$ (29)
with the incoming wave packets essentially re-emerging as outgoing packets with the same particle content and minimal radiation. The deeply broken point [Fig. 9(d), $\mu^{2}=-0.5$] exhibits similar qualitatively elastic behavior, $P_{11\rightarrow 11}\simeq 1$ and $\Delta t_{\mathrm{br2}}\simeq-177.781$, with even more clearly separated outgoing wave fronts due to the larger gap.
Near the critical coupling [Fig. 9(b)], the sandwich evolution ceases to produce a well-defined scattering event. Instead of the characteristic “X”-shaped collision pattern, $\langle\phi_{n}(t)\rangle$ develops a long-wavelength drift across the entire window, and no clean separation into incoming and outgoing wave packets is observed. This is not a numerical artifact or an instability of TDVP/uMPS, but a direct consequence of criticality: the sandwich construction assumes a finite correlation length $\ell$ that separates a central scattering region from asymptotic bulk domains. At the critical point the exact ground state is gapless ($m=0$) and scale invariant, with a diverging correlation length ($\ell\to\infty$), so any finite-window implementation of this protocol inevitably violates that assumption and fails to reach asymptotic scattering states. In this sense, the breakdown of the sandwich geometry at $\mu^{2}\approx\mu_{c}^{2}$ reflects the underlying physics of an infinitely correlated critical state rather than a flaw of the method, and it provides a dynamical signature of the quantum critical point that is complementary to the static FES diagnostics discussed above. Closer and closer to $\mu_{c}^{2}$ the required window sizes and evolution times grow with $\ell$, so in practice the protocol becomes rapidly more expensive and ultimately impractical as the critical point is approached.
@@ IV Summary and Discussion
We have extended TDVP-based uniform matrix product state techniques, previously applied to Ising field theory [36], to the interacting $\phi^{4}$ quantum field theory in $(1+1)$ dimensions. A finite-entanglement scaling analysis at $\lambda=0.8$ yields a controlled estimate of the critical mass-squared, $\mu_{c}^{2}\in]-0.2595,-0.2594[$, consistent with Ising universality and with the expected emergence of a gapless, conformal regime. This provides a quantitative map of the symmetric, near-critical, weakly broken, and deeply broken phases, anchoring the subsequent real-time simulations to a data-driven bracket of the critical point rather than to ad hoc parameter choices.
Using these uMPS ground states as asymptotic vacua, we simulated two-particle collisions in a sandwich geometry across the phase diagram and extracted both the elastic scattering probability $P_{11\to 11}(E)$ and the Wigner time delay $\Delta t(E)$. In the symmetric phase the collisions are strongly inelastic, with $P_{11\to 11}\simeq 0.712$, reflecting the sizable phase space for particle production in a nonintegrable theory with a fundamental boson. In the spontaneously broken phase the same protocol yields almost perfectly elastic collisions, $P_{11\to 11}\approx 1$, indicating that the excitations behave as stable, weakly radiating quasiparticles. As the system is tuned deeper into the broken regime, the increasing gap sharpens this behavior and further suppresses inelastic channels.
Near the critical coupling $\mu^{2}=\mu_{c}^{2}$ the sandwich evolution fails to produce a clean scattering event: instead of a well-defined “X” pattern in $\langle\phi_{n}(t)\rangle$, we observe a long-wavelength drift across the entire window and no clear separation into incoming and outgoing packets. This breakdown is a direct manifestation of criticality: the scattering construction relies on a finite correlation length to isolate a central interaction region from asymptotic bulk domains, an assumption that is invalidated when the mass gap closes and $\xi\to\infty$. The failure of the sandwich protocol at $\mu^{2}_{c}$ thus provides a dynamical signature of the quantum critical point, complementary to the static FES diagnostics.
@@ Acknowledgments
W.C. gratefully acknowledges the support provided by Jens Eisert, including financial assistance from Freie Universität Berlin through a research fellowship funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the Research Unit FOR 2724 “Thermische Maschinen in der Quantenwelt” (Fonds 0420611402).